a cylindrical fishing reel has a moment of inertia of I=6.8x10 to the negative 4th kgm squared and a radius of 4.8cm a friction clutch in the reel exerts a restraining torque of .75Nm if a fish pulls on the line the fisherman gets a bite and the reel begins to spin with an angular accelleration of 56 rad/s squared (a)what is the force of the fish on the line (b)how much line unwinds in .8s
Answer #1

I = moment of inertia of the reel = 6.8 x 10-4 kgm2

r = radius = 4.8 cm = 0.048 m

f = torque applied by friction clutch = 0.75 Nm

F = force applied by the fish

= Angular acceleration = 56 rad/s2

Torque equation is given as

F r - f = I

F (0.048) - 0.75 = (6.8 x 10-4) (56)

F = 16.4 N

b)

a = linear acceleration

Linear acceleration is given as

a = r

a = (0.048) (56)

a = 2.7 m/s2

t = time taken = 0.8 s

L = length of line unwound

length is given as

L = (0.5) a t2

L = (0.5) (2.7) (0.8)2

L = 0.864 m